# Isometries of 3-dimensional semi-Riemannian Lie groups

### v1

@@ 2.1.3. Elliptic type
$A$ has one real and two complex conjugate eigenvalues.
In this case, it can be easily shown that the real eigenvalue is zero with timelike eigenvector and that the complex conjugate eigenvalues are in fact imaginary with spacelike eigenvectors. The exponential of $A$, $\mathrm{exp}(A$) acting as an isomorphism of $V$ preserves a timelike direction $\ell$ and preserves its orthogonal plane $\mathscr{P}$ which is then spacelike. Futhermore, $\mathrm{exp}(A)$ acts as a rotation on $\mathscr{P}$.
@@ 2.3. Isometric automorphisms
We note that, by Lie’s second theorem, when $G$ is simply connected, the groups $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$ are in one-to-one correspondence. Another interesting point is that, in dimension 3, homomogeneous semi-Riemannian manifolds are always reductive when the isotropy is connected, [FR06]. Nevertheless, even in the case where $G$ acts transitively on $L/I$, it is possible to have many reductive complements but none that is a subalgebra, see Subsec.8.1 for an example.
@@ 3. $L$ solvable with abelian derived Killing algebra
In this section, we will analyse the case where $\mathfrak{N}$ is abelian. The case where $\mathfrak{N}=\mathfrak{heis}$ will be treated in section 4.
@@ 3.1. $\mathfrak{N}$ isomorphic to $\mathbb{R}^{3}$
The Lie algebra $\mathfrak{g}$ of $G$ is a 3-dimensional Lie algebra that can be identified with a subalgebra of $\mathfrak{L}$. If $\mathfrak{g}=\mathbb{R}^{3}$, then $G$ is Minskowski spacetime and therefore a spaceform. Let us then assume that $\mathfrak{g}\neq\mathbb{R}^{3}$. Thus the intersection of $\mathfrak{g}$ with $\mathbb{R}^{3}$ is necessarily of dimension $2$, let us say, $\mathfrak{g}\cap\mathbb{R}^{3}=\mathrm{span}\{X,Y\}$. Also $\mathfrak{g}$ contains a vector of the form $T+u$, with $u\in\mathbb{R}^{3}$, but since $\mathbb{R}^{3}$ is abelian, we can assume that $T$ is chosen such that $\mathfrak{g}=\mathrm{span}\{T,X,Y\}$.
@@ 3.2.1. Case where $\mathfrak{N}\subset\mathfrak{g}$
$ad_{U}([T,V])=[ad_{U}T,V]+[T,ad_{U}V]$ (3.1)
For $V\in\mathfrak{N}$, we have that $[ad_{U}T,V]=0$, since $\mathrm{ad}_{U}T\in\mathfrak{N}$ and $\mathfrak{N}$ is abelian. Therefore, from (3.1), we can write
$(\mathrm{ad}_{U}\circ\mathrm{ad}_{T})(V)=(ad_{T}\circ ad_{U})(V)$
@@
00 0 $=q(T,A(T))=\alpha q(T,X)+\beta q(T,Y)$
00 0 $=q(T,A(X))+q(A(T),X)=\alpha q(X,X)+\beta q(Y,X)$
00 0 $=q(T,A(Y))+q(A(T),Y)=q(T,X)+\alpha q(X,Y)+\beta q(Y,Y)$
00 0 $=q(X,A(X))=0$
00 0 $=q(X,A(Y))+q(A(X),Y)=q(X,X)$
00 0 $=q(Y,A(Y))=q(Y,X)$
@@
are the only non-vanishing terms of $m$. For $\mathfrak{sol}$, the metric $m$ is flat and it is complete, as proved in [BM08], thus the dimension of the Killing algebra is 6. As for $\mathfrak{h}(1),\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ and $\mathfrak{h}(\lambda)$ the metric $m$ is of negative sectional curvature but in all three cases it is incomplete, [CFZ25], so indeed the Killing algebra has dimension 4.
@@ 3.2.2. Case where $\mathfrak{N}\not\subset\mathfrak{g}$
Let $U$ be a generator of $\mathfrak{I}$. Then $U\notin\mathfrak{g}$ and, also, $U\notin\mathfrak{N}$, otherwise the rank of $\mathrm{iso}(U)$ would be less than 2. We then have $U=T+aX+bY+cZ$ with $a\neq 0$ or $c\neq 0$. We will show that $\mathrm{iso}(U)$ cannot be of elliptic type. Let $\pi:\mathfrak{L}\longrightarrow\mathfrak{L}/\mathfrak{I}$ be the canonical projection. For a given $V\in\mathfrak{L}$ we will write $\overline{V}$ for $\pi(V)$. Then $\{\overline{X},\overline{Y},\overline{Z}\}$ form a basis of $\mathfrak{L}/\mathfrak{I}$. Suppose that $[T,X]=\alpha Y+\beta T$ and $[T,Z]=\gamma Y+\delta T$, for some constants $\alpha,\beta,\gamma,\delta\in\mathbb{R}$. In the basis $\{\overline{Y},\overline{Z},\overline{X}\}$, $\mathrm{iso}(U)$ is given by the matrix
@@
This metric is flat but it is incomplete, [CFZ25], therefore the dimension of the isometry Lie algebra cannot be equal to $6$ and so it is either 3 or 4.
Considering $X^{\prime}=X-\tfrac{\alpha}{\gamma}Z$, the basis $\{X^{\prime},Y,Z\}$ spans $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ and the non-zero bracket relations in $\mathfrak{L}$ are $[X^{\prime},Y]=Y$, $[X^{\prime},T]=T,[T,Z]=\gamma Y$. Therefore $\mathfrak{L}$ can also be written as $\mathfrak{aff}(\mathbb{R})\ltimes\mathbb{R}^{2}$. We encounter here the case (5) of Th. 1.1. in [ABZ23] which is isometric to half-Minkowski space. Thus, the action of $\mathrm{Aff}^{+}(\mathbb{R})\times\mathbb{R}$ on $L/I$ is transitive (see Sec. 7 for more details).
@@
Let $\mathfrak{g}=\mathrm{span}\{X,Y,Z\}$ with $[X,Y]=Z$. Consider another element $T\in\mathfrak{L}$ such that $\mathfrak{N}=\mathrm{span}\{T,Z\}$. Since $\mathfrak{N}$ is abelian, $[Z,T]=0$. Thus, $Z$ is central in $\mathfrak{L}$. Let $U$ be a generator of the isotropy $\mathfrak{I}$. Then, up to scaling, $U=T+aX+bY+cZ$. Suppose that $[X,T]=\alpha Z+\beta T$ and $[T,Z]=\gamma Z+\delta T$, for some constants $\alpha,\beta,\gamma,\delta\in\mathbb{R}$, with $\beta\neq 0$ or $\delta\neq 0$. The Jacobi identity for $\mathfrak{L}$ implies that $\alpha\beta-\delta\gamma=0$. In the basis $\{\overline{Z},\overline{X},\overline{Y}\}$, $\mathrm{iso}(U)$ is given by the matrix
@@
Unsurprisingly, $\overline{Z}$ is a 0-eigenvector of $\mathrm{iso}(U)$. We observe that since the lower $2\times 2$-block matrix has zero determinant and its trace is zero, then $\lambda=0$ is the only eigenvalue of $\mathrm{iso}(U)$. Therefore, $iso(U)$ is necessarily nilpotent. Consider $V=\delta X-\beta Y$. We have that $V\neq 0$ and $[T,V]=0$. Moreover, $[U,V]=-(b\delta+a\beta)Z=0$. Hence, $V$ is another 0-eigenvector of $\mathrm{iso}(U)$, which cannot happen for a non-trivial Lorentzian infinitesimal isometry. Therefore, this case cannot occur.
3.3. $\mathfrak{N}$ is isomorphic to $\mathbb{R}$
@@ 4.2. Case where $\mathfrak{g}\neq\mathfrak{heis}$
Consider the isotropy $\mathrm{iso}(X)$ on $\mathfrak{L}/\mathbb{R}X$. As usual, we can consider the basis $\{\overline{T+cX},\overline{Y},\overline{Z}\}$ of $\mathfrak{L}/I$ given by the projection $\mathfrak{L}\longrightarrow\mathfrak{L}/I$. Given $\alpha,\gamma,\delta\in\mathbb{R}$ such that $[T,X]=\alpha X+\gamma Y+\delta Z$, we have
$\begin{array}[]{rl}\mathrm{iso}(X)(T+cX)&=\,\overline{\mathrm{ad}_{X}(T)}={-\overline{(\alpha X+\gamma Y+\delta Z)}}=-\gamma\overline{Y}-\delta\overline{Z},\\ \mathrm{iso}(X)(\overline{Y})&=\,\overline{\mathrm{ad}_{X}(Y)}=\overline{Z},\\ \mathrm{iso}(X)(\overline{Z})&=\,\overline{\mathrm{ad}_{X}(Z)}=0.\\ \end{array}$
@@
Therefore, $\mathrm{iso}(X)$ must be 3-step nilpotent. If so, then $[T,X]$ has non-trivial $Y$ component, i.e. $\gamma\neq 0$. Since $[\mathfrak{L},\mathfrak{L}]=\mathfrak{heis}$ and $\mathfrak{g}$ is a subalgebra, then $[\mathfrak{g},\mathfrak{g}]\subset\mathfrak{n}$. In particular, there exists some $\beta,b\in\mathbb{R}$ such that $[T,Y]=\beta Y+bZ$. In sum, and recalling that $\mathrm{ad}_{T}|_{\mathfrak{heis}}$ is a derivation of $\mathfrak{heis}$, then $\mathrm{ad}_{T}|_{\mathfrak{heis}}$ is represented by the matrix
@@
Notice that $\alpha\neq 0$ and $\beta\neq 0$ are given by the fact that $[\mathfrak{L},\mathfrak{L}]=\mathfrak{heis}$. Observe, furthermore, that this implies that $\mathfrak{g}$ cannot be an ideal of $\mathfrak{L}$.
@@ 4.3. The Lie algebra $\mathfrak{g}$
$[S,Y]=\beta Y+(b+c)Z,\quad[S,Z]=(\alpha+\beta)Z,\quad[Y,Z]=0.$
@@
If $\alpha+\beta\neq 0$ then, $\mathfrak{g}_{B}$ is isomorphic to one of $\mathfrak{h}(\lambda)$, $10<|\lambda|<$ 1, or to the unimodular $\mathfrak{sol}$. Notice that, up to replacing $Y$ by $\tilde{Y}=Y-\frac{b}{\alpha+\beta}Z$, we can assume that $b=0$ in the matrix $A$. If $\alpha+\beta=0$, then $\mathfrak{g}_{B}$ is isomorphic to $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$.
@@ 4.4.2. Case 2: $\mathfrak{g}\neq\mathfrak{heis}$ with $\alpha+\beta\neq 0$
We remark that when computing the orbit of $q$ under the action of the automorphism group of $\mathfrak{h}(\lambda)$, $10<|\lambda|<$ 1, we have to take into consideration that either $\overline{Y}$ or $\overline{Z}$ can take on the role of $e_{3}$ as given in Table 1. An exception is made for the case of the Lie algebra $\mathfrak{sol}$ where these two metrics are, in fact, equivalent by means of the extra component in the automorphism group.
@@ 4.4.3. Case 3: $\mathfrak{g}\neq\mathfrak{heis}$ with $\alpha+\beta=0$
In this case, $\mathfrak{g}$ is $\mathrm{aff}(\mathbb{R})\oplus\mathbb{R}$. The center of $\mathfrak{g}$ is, thus, generated by $Z$, the center of $\mathfrak{heis}$. We remark that $\alpha$ and $\beta$ do not contribute to the components of metric associated to the the nilpotent infinitesimal isometry and that the automorphism group of $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ is the same as the automorphism group of $\mathfrak{h}(\lambda)$, $10<|\lambda|<$ 1. Therefore, the metrics in question are equivalent to the metric with normal form $m_{1}$.
@@ 4.5. Space of plane waves
We take $X^{\prime}=X$. For $Y$ we rescale it by some $r\neq 0$ and perturb it by $X$, hence, $Y\mapsto Y^{\prime}=sX+rY$ (we can also perturb it by $Z$, but this is enough for our aim). Since we are only changing the basis by $\mathrm{Aut}(\mathfrak{heis})$, we have no freedom in $Z$, it automatically changed to $Z^{\prime}=rZ$. The brackets of $\widetilde{T}$ in the new basis are given by

### v2

@@ 2.1.3. Elliptic type
AA has one real and two complex conjugate eigenvalues.
In this case, it can be easily shown that the real eigenvalue is zero with timelike eigenvector and that the complex conjugate eigenvalues are in fact imaginary with spacelike eigenvectors. The exponential of $A$, $\mathrm{exp}(A$) acting as an isomorphism of $V$ preserves a timelike direction $\ell$ and preserves its orthogonal plane $\mathscr{P}$ which is then spacelike. Furthermore, $\mathrm{exp}(A)$ acts as a rotation on $\mathscr{P}$.
@@ 2.3. Isometric automorphisms
We note that, by Lie’s second theorem, when $G$ is simply connected, the groups $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$ are in one-to-one correspondence. Another interesting point is that, in dimension 3, homogeneous semi-Riemannian manifolds are always reductive when the isotropy is connected, [FR06]. Nevertheless, even in the case where $G$ acts transitively on $L/I$, it is possible to have many reductive complements but none that is a subalgebra, see Subsec.8.1 for an example.
@@ 3. $L$ solvable with abelian derived Killing algebra
In this section, we will analyze the case where $\mathfrak{N}$ is abelian. The case where $\mathfrak{N}=\mathfrak{heis}$ will be treated in section 4.
@@ 3.1. $\mathfrak{N}$ isomorphic to $\mathbb{R}^{3}$
The Lie algebra $\mathfrak{g}$ of $G$ is a 3-dimensional Lie algebra that can be identified with a subalgebra of $\mathfrak{L}$. If $\mathfrak{g}=\mathbb{R}^{3}$, then $G$ is Minkowski spacetime and therefore a space form. Let us then assume that $\mathfrak{g}\neq\mathbb{R}^{3}$. Thus the intersection of $\mathfrak{g}$ with $\mathbb{R}^{3}$ is necessarily of dimension $2$, let us say, $\mathfrak{g}\cap\mathbb{R}^{3}=\mathrm{span}\{X,Y\}$. Also $\mathfrak{g}$ contains a vector of the form $T+u$, with $u\in\mathbb{R}^{3}$, but since $\mathbb{R}^{3}$ is abelian, we can assume that $T$ is chosen such that $\mathfrak{g}=\mathrm{span}\{T,X,Y\}$.
@@ 3.2.1. Case where $\mathfrak{N}\subset\mathfrak{g}$
$\mathrm{ad}_{U}([T,V])=[\mathrm{ad}_{U}T,V]+[T,\mathrm{ad}_{U}V]$ (3.1)
For $V\in\mathfrak{N}$, we have that $\mathrm{ad}_{U}T,V]=0$, since $\mathrm{ad}_{U}T\in\mathfrak{N}$ and $\mathfrak{N}$ is abelian. Therefore, from (3.1), we can write
$(\mathrm{ad}_{U}\circ\mathrm{ad}_{T})(V)=(\mathrm{ad}_{T}\circ\mathrm{ad}_{U})(V)$
@@
$0$ $=q(T,A(T))=\alpha q(T,X)+\beta q(T,Y)$
$0$ $=q(T,A(X))+q(A(T),X)=\alpha q(X,X)+\beta q(Y,X)$
$0$ $=q(T,A(Y))+q(A(T),Y)=q(T,X)+\alpha q(X,Y)+\beta q(Y,Y)$
$0$ $=q(X,A(X))=0$
$0$ $=q(X,A(Y))+q(A(X),Y)=q(X,X)$
$0$ $=q(Y,A(Y))=q(Y,X)$
@@
are the only non-vanishing terms of $m$. For $\mathfrak{sol}$, the metric $m$ is flat and it is complete, as proved in [BM08], thus the dimension of the Killing algebra is 6. As for $\mathfrak{h}(1),\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ and $\mathfrak{h}(\lambda)$ the metric $m$ is of negative sectional curvature but in all three cases it is incomplete, [CF25], so indeed the Killing algebra has dimension 4.
@@ 3.2.2. Case where $\mathfrak{N}\not\subset\mathfrak{g}$
Let $U$ be a generator of $\mathfrak{I}$. Then $U\notin\mathfrak{g}$ and, also, $U\notin\mathfrak{N}$, otherwise the rank of $\mathrm{iso}(U)$ would be less than 2. We then have $U=T+aX+bY+cZ$ with $a\neq 0$ or $c\neq 0$. We will show that $\mathrm{iso}(U)$ cannot be of elliptic type. Let $\pi:\mathfrak{L}\longrightarrow\mathfrak{L}/\mathfrak{I}$ be the canonical projection. For a given $V\in\mathfrak{L}$ we will write $\overline{V}$ for $\pi(V)$. Then $\{\overline{X},\overline{Y},\overline{Z}\}$ form a basis of $\mathfrak{L}/\mathfrak{I}$. Suppose that $T,X]=\alpha Y+\beta T$ and $T,Z]=\gamma Y+\delta T$, for some constants $\alpha,\beta,\gamma,\delta\in\mathbb{R}$. In the basis $\{\overline{Y},\overline{Z},\overline{X}\}$, $\mathrm{iso}(U)$ is given by the matrix
@@
This metric is flat but it is incomplete, [CF25], therefore the dimension of the isometry Lie algebra cannot be equal to $6$ and so it is either 3 or 4.
Considering $X^{\prime}=X-\tfrac{\alpha}{\gamma}Z$, the basis $\{X^{\prime},Y,Z\}$ spans $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ and the non-zero bracket relations in $\mathfrak{L}$ are $X^{\prime},Y]=Y$, $X^{\prime},T]=T,[T,Z]=\gamma Y$. Therefore $\mathfrak{L}$ can also be written as $\mathfrak{aff}(\mathbb{R})\ltimes\mathbb{R}^{2}$. We encounter here the case (5) of Th. 1.1. in [ABZ23] which is isometric to half-Minkowski space. Thus, the action of $\mathrm{Aff}^{+}(\mathbb{R})\times\mathbb{R}$ on $L/I$ is transitive (see Sec. 7 for more details).
@@
Let $\mathfrak{g}=\mathrm{span}\{X,Y,Z\}$ with $[X,Y]=Z$. Consider another element $T\in\mathfrak{L}$ such that $\mathfrak{N}=\mathrm{span}\{T,Z\}$. Since $\mathfrak{N}$ is abelian, $[Z,T]=0$. Thus, $Z$ is central in $\mathfrak{L}$. Let $U$ be a generator of the isotropy $\mathfrak{I}$. Then, up to scaling, $U=T+aX+bY+cZ$. Suppose that $X,T]=\alpha Z+\beta T$ and $T,Z]=\gamma Z+\delta T$, for some constants $\alpha,\beta,\gamma,\delta\in\mathbb{R}$, with $\beta\neq 0$ or $\delta\neq 0$. The Jacobi identity for $\mathfrak{L}$ implies that $\alpha\beta-\delta\gamma=0$. In the basis $\{\overline{Z},\overline{X},\overline{Y}\}$, $\mathrm{iso}(U)$ is given by the matrix
@@
Unsurprisingly, $\overline{Z}$ is a 0-eigenvector of $\mathrm{iso}(U)$. We observe that since the lower $2\times 2$-block matrix has zero determinant and its trace is zero, then $\lambda=0$ is the only eigenvalue of $\mathrm{iso}(U)$. Therefore, $\mathrm{iso}(U)$ is necessarily nilpotent. Consider $V=\delta X-\beta Y$. We have that $V\neq 0$ and $[T,V]=0$. Moreover, $U,V]=-(b\delta+a\beta)Z=0$. Hence, $V$ is another 0-eigenvector of $\mathrm{iso}(U)$, which cannot happen for a non-trivial Lorentzian infinitesimal isometry. Therefore, this case cannot occur.
$3.$ $\mathfrak{N}$ is isomorphic to $\mathbb{R}$
@@ 4.2. Case where $\mathfrak{g}\neq\mathfrak{heis}$
Consider the isotropy $\mathrm{iso}(X)$ on $\mathfrak{L}/\mathbb{R}X$. As usual, we can consider the basis $\{\overline{T+cX},\overline{Y},\overline{Z}\}$ of $\mathfrak{L}/I$ given by the projection $\mathfrak{L}\longrightarrow\mathfrak{L}/I$. Given $\alpha,\gamma,\delta\in\mathbb{R}$ such that $T,X]=\alpha X+\gamma Y+\delta Z$, we have
[…]
@@
Therefore, $\mathrm{iso}(X)$ must be 3-step nilpotent. If so, then $[T,X]$ has non-trivial $Y$ component, i.e. $\gamma\neq 0$. Since $\mathfrak{L},\mathfrak{L}]=\mathfrak{heis}$ and $\mathfrak{g}$ is a subalgebra, then $\mathfrak{g},\mathfrak{g}]\subset\mathfrak{n}$. In particular, there exists some $\beta,b\in\mathbb{R}$ such that $T,Y]=\beta Y+bZ$. In sum, and recalling that $\mathrm{ad}_{T}|_{\mathfrak{heis}}$ is a derivation of $\mathfrak{heis}$, then $\mathrm{ad}_{T}|_{\mathfrak{heis}}$ is represented by the matrix
@@
Notice that $\alpha\neq 0$ and $\beta\neq 0$ are given by the fact that $\mathfrak{L},\mathfrak{L}]=\mathfrak{heis}$. Observe, furthermore, that this implies that $\mathfrak{g}$ cannot be an ideal of $\mathfrak{L}$.
@@ 4.3. The Lie algebra $\mathfrak{g}$
$S,Y]=\beta Y+(b+c)Z,\quad[S,Z]=(\alpha+\beta)Z,\quad[Y,Z]=0.$
@@
If $\alpha+\beta\neq 0$ then, $\mathfrak{g}_{B}$ is isomorphic to one of $\mathfrak{h}(\lambda)$, $0<|\lambda|<1$, or to the unimodular $\mathfrak{sol}$. Notice that, up to replacing $Y$ by $\tilde{Y}=Y-\frac{b}{\alpha+\beta}Z$, we can assume that $b=0$ in the matrix $A$. If $\alpha+\beta=0$, then $\mathfrak{g}_{B}$ is isomorphic to $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$.
@@ 4.4.2. Case 2: $\mathfrak{g}\neq\mathfrak{heis}$ with $\alpha+\beta\neq 0$
We remark that when computing the orbit of $q$ under the action of the automorphism group of $\mathfrak{h}(\lambda)$, $0<|\lambda|<1$, we have to take into consideration that either $\overline{Y}$ or $\overline{Z}$ can take on the role of $e_{3}$ as given in Table 1. An exception is made for the case of the Lie algebra $\mathfrak{sol}$ where these two metrics are, in fact, equivalent by means of the extra component in the automorphism group.
@@ 4.4.3. Case 3: $\mathfrak{g}\neq\mathfrak{heis}$ with $\alpha+\beta=0$
In this case, $\mathfrak{g}$ is $\mathrm{aff}(\mathbb{R})\oplus\mathbb{R}$. The center of $\mathfrak{g}$ is, thus, generated by $Z$, the center of $\mathfrak{heis}$. We remark that $\alpha$ and $\beta$ do not contribute to the components of metric associated to the nilpotent infinitesimal isometry and that the automorphism group of $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ is the same as the automorphism group of $\mathfrak{h}(\lambda)$, $0<|\lambda|<1$. Therefore, the metrics in question are equivalent to the metric with normal form $m_{1}$.
@@ 4.5. Space of plane waves
We take $X^{\prime}=X$. For $Y$ we rescale it by some $r\neq 0$ and perturb it by $X$, hence, $Y\mapsto Y^{\prime}=sX+rY$ (we can also perturb it by $Z$, but this is enough for our aim). Since we are only changing the basis by $\mathrm{Aut}(\mathfrak{heis})$, we have no freedom in $Z$, it automatically changes to $Z^{\prime}=rZ$. The brackets of $\widetilde{T}$ in the new basis are given by